grades on midterm exam given with mean 75 and std dev 9.F grade is all scores below 60th percentile of distribution.find the lowest acceptable score.
Answer by stanbon(75887)

grades on midterm exam given with mean 75 and std dev 9.
F grade is all scores below 60th percentile of distribution.
find the lowest acceptable score.
---
Find the z-value that has a 0.60 left tail:
invNorm(0.6) = 0.2533
-----
Use x = zs+u to find the corresponding exam score:
x = 0.2533*9+75 = 77.28
==========================
Cheers,
Stan H.